# VQA2: Visual Question Answering for Video Quality Assessment

### v2

@@ Abstract
footnotetext: *Equal contribution. Corresponding author.
@@ 5.2 Evaluation on Quality Scoring Tasks
frames for motion extraction resolution $224\times 224$ /
@@ Appendix A Detailed Information About Training
Model Structure and Training Hyperparameters.
@@
The training of our Assistant model is based on the UGC-Scorer, however bypassing the Stage-2 streaming video data. Additionally, we design distinct specialized Scorers for typical UGC (offline) videos and streaming videos (featuring long-time stalling and rebuffering) instead of using a unified model. The reasons for this design are as follows.
@@ B.1 Annotation Preparation
Table 11: The Common Quality Attributes Examples. The quality attributes in bold are those that must be considered in most videos. The quality attributes displayed in italic font are the ones that can only be noticed in a small number of videos. The remaining attributes that are not marked should be considered depending on the content and quality level of the specific video.
@@
Table 12: Degree descriptions and their criterion. To minimize the annotation bias, we provide the specific applicable range and criterion for each degree description. (Given the unavoidable perceptual biases that exist with individual differences, the provided applicable ranges are intentionally designed to be flexible.)
@@ B.2 Overall Quality Depiction Settings
  • The overall quality depiction should clearly state the spatial area (the whole picture or a specific area) corresponding to each quality attribute under the temporal description. If a quality attribute exists in the whole frame, such a location description can be omitted.
@@
  • Good (60-80) This part of the videos gives good overall impressions, but most have noticeable flaws. The depiction process needs to consider both positive and negative quality attributes. The annotator should try to use a transitional tone, which is highlighted in red below. (Example: The video has good sharpness throughout the playback, and the scene can be distinguished to a high degree. However, there are flickers with high frequency and minor amplitude during the video playback process, and the light is relatively dark and unnatural when the camera switches to the woman in the middle, which affects the viewer’s viewing experience to a certain degree.)
  • Fair (40-60) There are obvious flaws in this part of the video, and negative quality attributes (distortions) have become the main factor affecting the evaluation basis. The overall description only needs to describe the distortion present in the video. (Example: the video has poor clarity throughout the playback process, there are high-frequency, irregular, and high-amplitude camera shakes, and the color richness of in the second half of the video is low. These distortions seriously affect the viewer’s viewing experience.) A detailed description of this distortion can be added when necessary, such as when the video has only a dominant distortion. (Example: There is relatively severe compression blur throughout the video playback, which makes the faces of the main characters in the video barely distinguishable, and the details are seriously lost.)
  • Poor (20-40) This part of the video has more serious distortions. The annotation requirements are almost consistent with the Fair level. (Example: The video has very poor sharpness throughout the playback. There are high-frequency, irregular and obvious lens shakes, and the color richness during $3$-$7$ seconds is very low. These distortions seriously affect the viewer’s viewing experience.)
@@ B.3 In-context Depiction/Extended Conversation Settings
If the annotator chooses the in-context depiction, they are required to find one or more representative time points in the video/specific locations in the picture and describe their quality. If they need to describe a specific moment, the in-context temporal description shown in Tab. 13 is preferable.
@@
  • Design a question and response focusing on the quality of a particular moment, period or certain spatial location within the video (similar to the in-context depiction rephrased in an conversational format).
@@ B.4 Special Designs for AIGC-video Annotations
Since AIGC video quality evaluation focuses more on visual realism, factual consistency, and motion coherence—rather than solely on low-level visual quality—our approach to creating overall quality depictions for the $998$ AIGC videos in the dataset differs from the previous requirements. Specifically, we instruct annotators to provide detailed descriptions across $5$ aspects provided in the original dataset: visual quality, temporal alignment, dynamic degree, text-to-video alignment, and factual consistency. We use the quality levels (ranging from $1$ to $4$) annotated in the dataset as references for overall depictions. Additionally, given the fact that the length of these videos is generally very short (typically less than $5$ seconds with a low frame rate), the temporal description is not necessary. Additionally, we bypass the in-context depiction / extended conversation annotation step. The GPT rewriting process is also solely based on the $5$ dimensions in the overall depiction.
@@ B.5 Overview of Subjective Experiment Setup
The subjective annotation experiment involves $33$ participants, all well-educated undergraduate or graduate students with relevant professional backgrounds. Among them, $16$ participants are experts in related fields. The selection of participants ensures the quality of the human-annotated data. The entire experiment lasts one month, during which all participants are required to view and annotate videos in environments meeting the requirements of formal subjective experiments. The number of videos annotated by each participant ranged from $100$ to $1,000$. To prevent fatigue, participants are instructed to evenly distribute their daily annotation tasks, with each individual annotating no less than $15$ videos and no more than $30$ videos per day.
@@ C.1 Prompts design for GPT-extension
You will receive a detailed depiction of the quality of a video, which may include evaluations and comments on various attributes of the video’s quality. Based on this description, create exactly $3$ questions about specific video quality attributes (sharpness, fluency, flicker, etc.) and provide the corresponding answers. The questions and answers can be in the form of yes/no choices (e.g., Question: Is the sharpness of the video high throughout the playback? A. Yes B. No Answer: A.), multi-choices (single answer) questions (e.g., Question: How is the degree of the sharpness of the video throughout the playback? A. High B. Relatively good C. Relatively poor D. Extremely low. Answer: B.) or open-ended responses (e.g., Question: How would you rate the sharpness of the video throughout playback? Answer: Average/Excellent/Poor). Please ensure that the questions and answers are entirely based on the video quality depiction without including any information beyond it. Below is an example:
@@
Please note that questions like “How is the brightness of the video?’ are not suitable as there is no absolute description of brightness quality in the description, only a mention of its decrease. Please format the output as follows: Question 1: Question: Answer: Question 2: Question: Answer: Question 3: Question: Answer:. You need to simulate as if you have seen the video; therefore, all questions and answers should not include any traces of the provided text depiction.
@@
    Temporal analysis, which asks about and answers the relationship between the quality of two different time points or segments within the video.
@@
    Temporal retrieval, which provides a description of a quality or distortion event and then asks when this event occurs in the video or how frequently it appears.
@@
Please note that you will not see the video itself but only the text depiction provided, so all questions and answers must strictly adhere to the description. If the description does not include relevant information, such information should not appear in the questions or answers you design. You may design questions through single-answer choices or open-ended responses. Please format the output as follows: Question: Answer:. The video quality depiction is: [overall depiction]. You need to simulate as if you have seen the video; therefore, all questions and answers should not include any traces of the provided text depiction.
@@
    Example: Depiction: The video has low brightness and faded colors. Question: What adjustments could be made to re-shoot the video to achieve higher quality? Answer: Shooting on a sunnier day or increasing the overall brightness of the surrounding shooting environment would improve quality.
Please note that you will not see the video; only the textual depiction will be provided. Therefore, all questions and answers must strictly adhere to the content of the depiction. If the depiction explicitly states the cause of a quality phenomenon, questions of type 1 must strictly follow the reason given in the depiction. If no explicit cause is stated in the depiction, do not design type 1 questions. Please format the output as follows: Question: Answer:. The video quality depiction is: [overall depiction]. You need to simulate as if you have seen the video; therefore, all questions and answers should not include any traces of the provided text depiction.
@@
You will receive one brief in-context quality description, which primarily describes a positive or negative quality phenomenon occurring in a specific moment or period of a video or space within the frames. Based on this description, please rewrite only one question. Requirements: If the in-context quality description targets a specific moment or period in the video or a particular object/part of the video space (or a combination of both), the rewritten question must focus on that specific time or spatial part/object stated in the description.
Example: Depiction: Out-of-focus blur occurs when the person starts running in the video. Question: At what moment does defocus blur occur in the video? Please describe this moment precisely. Answer: When the person starts running in the video.)
@@ C.2 System Prompts Design
A specific system prompt is added before the instruction Q&A pairs to facilitate the LMM’s understanding of the exact meaning of the question and its corresponding answer.
@@
When evaluating quality scoring tasks, we choose a relatively simple system prompts design to guarantee that the evaluation process is end-to-end (without any additional temporal information that may need manual extraction like length, frame rate, and stalling information). The system prompt design is shown below.
@@ Appendix D Quality Scoring Method
where quality_levels denotes the list of transformed quality levels: [High, good, fair, poor, low] and $\mathcal{P}$ represents the model’s output probability of each quality level. Specifically, we first extract the vector (with its dimension matching the tokenizer vocabulary size) at the model’s output sequence corresponding to the quality-level description word (the $-3$ index in our model). We then take the values at the indexes in this vector associated with each quality level in the tokenizer’s vocabulary (indexes $1550$, $1661$, $6624$, $7852$, and $3347$ in our model). Then, we apply softmax normalization to these $5$ values. $\omega$ denotes the weights assigned to the normalized probabilities of these $5$ quality levels ($[1,0.75,0.5,0.25,0]$). The weighted sum of these probabilities yields the predicted quality score $\mathcal{Q}$, ranging between $[0,1]$.
@@
Figure 9: A video with extremely high color temperature, very low sharpness, and frequent, severe stuttering. The case demonstrates that the Assistant model can accurately identify videos with low fluency, perceive the level of color temperature, give reasonable suggestions for quality refinement, and provide a relatively accurate scoring for the low-quality videos.
Figure 10: A video shot on a mobile phone with a stable camera trajectory. The video undergoes some compression during file transfer, slightly reducing its sharpness. The beginning shows noticeable overexposure on the window, which does not persist throughout, resulting in overall medium quality. This example demonstrates that the Assistant model can accurately detect camera shake, perform precise spatial-temporal retrieval or analysis of distortions, and provide a reasonably accurate rating for a medium-quality video.
@@ Appendix F Applications
We created a demo based on Gradio to test the application value of the Assistant model in video quality question-answering. Below, we provide four specific examples in Fig. 9, 10, 11, and 12) .
@@ Appendix G Limitations
Due to time constraints, we can not conduct extensive experiments to obtain a larger amount of distorted recognition pretraining data. We utilize the data from existing studies and some synthetically generated data. Therefore, the Stage-1 can be regarded as a ”task-specific pretraining for video quality assessment” on small-scale data rather than the large-scale pretraining typically used to construct foundation models in comprehensive LMMs.
@@
Despite employing various methods, such as GPT-based extension and refinement and quality-label-referenced attribute selection to minimize annotation biases, the individual perception biases can still not be entirely eliminated. This limitation slightly impacts the diversity of model outputs and the accuracy of responses to specific questions.
@@ Appendix H Acknowledgements
In the subjective annotation experiments, all participants are thoroughly informed in advance about the workload. Participation in the experiment is entirely voluntary without any enforcement. Throughout the experiment, no participants report experiencing fatigue or discomfort. Upon completion of the experiment, we provide all participants with appropriate compensation. We would like to express our heartfelt gratitude to all participants for their contributions to the subjective experiments.
@@ Appendix I License
All outputs of our work, including the complete data of the VQA2 Instruction Dataset as well as the code and fine-tuned weights of the VQA2 Series Models, will be fully open-sourced and made freely available to the research community for further studies.

### v3

@@ Abstract
footnotetext: *Equal contribution. Project leader. Corresponding author.
@@ 5.2 Evaluation on Quality Scoring Tasks
Frames (for motion extraction) Resolution $224\times 224$ /
@@ Appendix A Detailed Information About Training
Model Structure and Training Hyper-parameters.
@@
As described in 4.3, the training of our Assistant model is based on the UGC-Scorer, however bypassing the Stage-2 streaming video data. Additionally, we design distinct specialized Scorers for typical UGC (offline) videos and streaming videos (featuring long-time stalling and rebuffering) instead of using a unified model. The reasons for this design are as follows.
@@ B.1 Annotation Preparation
Table 11: The representative quality attributes examples. The quality attributes in bold are those that must be considered in most videos. The quality attributes displayed in italic font are the ones that can only be noticed in a small number of videos. The remaining attributes that are not marked should be considered depending on the content and quality level of the specific video.
@@
Table 12: Degree descriptions and their criterion. To minimize the annotation bias, we provide the specific applicable range and criterion for each degree description level. (Given the unavoidable perceptual biases that exist with individual differences, the provided applicable ranges are intentionally designed to be flexible.)
@@ B.2 Overall Quality Depiction Settings
  • The overall quality depiction should clearly state the spatial area (the whole picture or a specific area) corresponding to each quality attribute under the temporal description. If a quality attribute exists in the whole frame (such as low sharpness caused by low resolution), such a location description can be omitted.
@@
  • Good (60-80) This part of the videos gives good overall impressions, but most have noticeable flaws. The depiction process needs to consider both positive and negative quality attributes. The annotator should try to use a transitional tone, which is highlighted in red below (Example: The video has good sharpness throughout the playback, and the scene can be distinguished to a high degree. However, there are flickers with high frequency and minor amplitude during the video playback process, and the light is relatively dark and unnatural when the camera switches to the woman in the middle, which affects the viewer’s viewing experience to a certain degree).
  • Fair (40-60) There are obvious flaws in this part of the video, and negative quality attributes (distortions) have become the main factor affecting the evaluation basis. The overall depiction only needs to describe the distortion present in the video (Example: the video has poor clarity throughout the playback process, there are high-frequency, irregular, and high-amplitude camera shakes, and the color richness of in the second half of the video is low. These distortions seriously affect the viewer’s viewing experience.) When the video has only a dominant distortion, a detailed description of this distortion can be added for extended depiction. (Example: There is relatively severe compression blur throughout the video playback, which makes the faces of the main characters in the video barely distinguishable, and the edge details are seriously lost).
  • Poor (20-40) This part of the video has more serious distortions. The annotation requirements are almost consistent with the Fair level (Example: The video has very poor sharpness throughout the playback. There are high-frequency, irregular and obvious lens shakes, and the color richness during $3$-$7$ seconds is very low. These distortions seriously affect the viewer’s viewing experience).
@@ B.3 In-context Depiction/Extended Conversation Settings
As mentioned in 3.4, if the annotator chooses the in-context depiction, they are required to find one or more representative time points in the video/specific locations in the picture and describe their quality. If they need to describe a specific moment, the in-context temporal description shown in Tab. 13 is preferable.
@@
  • Design a question and response focusing on the quality of a particular moment, period, or certain spatial location within the video (similar to the in-context depiction rephrased in a conversational format).
@@ B.4 Special Designs for AIGC-video Annotations
Since AIGC video quality evaluation focuses more on visual realism, factual consistency, and motion coherence—rather than solely on low-level visual quality—our approach to creating overall quality depictions for the $998$ AIGC videos in the dataset differs from the previous requirements. Specifically, we instruct annotators to provide detailed descriptions across $5$ aspects provided in the original dataset [he2024videoscore]: visual quality, temporal alignment, dynamic degree, text-to-video alignment, and factual consistency. We use the quality levels (ranging from $1$ to $4$) annotated in the dataset as references for overall depictions. Additionally, given the fact that the length of these videos is generally very short (typically less than $5$ seconds with a low frame rate), the temporal description is not necessary. Additionally, we bypass the in-context depiction / extended conversation annotation step. The GPT rewriting process is also solely based on the $5$ dimensions in the overall depiction.
@@ B.5 Overview of Subjective Experiment Setup
The subjective annotation experiment involves $33$ participants, all well-educated undergraduate or graduate students with relevant professional backgrounds. Among them, $16$ participants are experts in related fields. The selection of participants ensures the quality of the human-annotated data. The entire experiment lasts one month, during which all participants are required to view and annotate videos in environments meeting the requirements of formal subjective experiments. The number of videos annotated by each participant ranges from $100$ to $1,000$. To prevent fatigue, participants are instructed to evenly distribute their daily annotation tasks, with no less than $15$ videos and no more than $30$ videos per day.
@@ C.1 Prompts design for GPT-extension
You will receive a detailed depiction of the quality of a video, which may include evaluations and comments on various attributes of the video’s quality. Based on this depiction, create exactly $3$ questions about specific video quality attributes (sharpness, fluency, flicker, etc.) and provide the corresponding answers. The questions and answers can be in the form of yes/no choices (e.g., Question: Is the sharpness of the video high throughout the playback? A. Yes B. No Answer: A.), multi-choices (single answer) questions (e.g., Question: How is the degree of the sharpness of the video throughout the playback? A. High B. Relatively good C. Relatively poor D. Extremely low. Answer: B.) or open-ended responses (e.g., Question: How would you rate the sharpness of the video throughout playback? Answer: Average/Excellent/Poor). Please ensure that the questions and answers are entirely based on the video quality depiction without including any information beyond it. Below is an example:
@@
Please note that questions like “How is the brightness of the video?’ are not suitable as there is no absolute description of brightness quality in the description, only a mention of its decrease. Additionally, you will not see the video; only the textual depiction will be provided. Therefore, all questions and answers must strictly adhere to the depiction, and no information beyond what is described should be included. Please format the output as follows: Question 1: Question: Answer: Question 2: Question: Answer: Question 3: Question: Answer:. You need to simulate as if you have seen the video; therefore, all questions and answers should not include any traces of the provided text depiction.
@@
    Temporal analysis, which is about and answers the relationship between the quality of two different time points or segments within the video.
@@
    Temporal retrieval, which is about when a quality event occurs in the video or how frequently it appears.
@@
Please note that you will not see the video itself but only the text depiction provided, so all questions and answers must strictly adhere to the depiction. If the description does not include relevant information, such information should not appear in the questions or answers you design. You may design questions through single-answer choices or open-ended responses. Please format the output as follows: Question: Answer:. The video quality depiction is: [overall depiction]. You need to simulate as if you have seen the video; therefore, all questions and answers should not include any traces of the provided text depiction.
@@
    Example: Depiction: The video has low brightness and faded colors. Question: What adjustments could be made to re-shoot the video to achieve higher quality? Answer: Reshooting the video on a sunnier day or increasing the overall brightness of the surrounding shooting environment would improve quality.
Please note that you will not see the video; only the textual depiction will be provided. Therefore, all questions and answers must strictly adhere to the content of the depiction. If the depiction explicitly states the cause of a quality event, questions of type $1$ must strictly follow the reason given in the depiction. If no explicit cause is stated in the depiction, do not design type $1$ questions. Please format the output as follows: Question: Answer:. The video quality depiction is: [overall depiction]. You need to simulate as if you have seen the video; therefore, all questions and answers should not include any traces of the provided text depiction.
@@
You will receive one brief in-context quality depiction, which primarily describes a positive or negative quality phenomenon occurring in a specific moment or period of a video or space within the frames. Based on this depiction, please rewrite only $1$ question. Requirements: If the in-context quality description targets a specific moment or period in the video or a particular object/part of the frame space (or a combination of both), the rewritten question must focus on that specific time or spatial part/object stated in the depiction.
Example 1: Depiction: Out-of-focus blur occurs when the person starts running in the video. Question: At what moment does defocus blur occur in the video? Please describe this moment precisely. Answer: When the person starts running in the video.)
Example 2: Depiction: There is local overexposure on the full-body of the performers in the center and on the right. Question: Where in the frame does overexposure occur? Answer: On the full-body of the performers in the center and on the right side.)
@@ C.2 System Prompts Design
As mentioned in LABEL:setup, a specific system prompt is added before the instruction Q&A pairs to facilitate the LMM’s understanding of the exact meaning of the question and its corresponding answer.
@@
When evaluating quality scoring tasks, we choose a relatively simple system prompts design to guarantee that the evaluation process is end-to-end (without any additional temporal information that may need manual extraction, like length, frame rate, and stalling information). The system prompt design is shown below.
@@ Appendix D Quality Scoring Method
where quality_levels denotes the list of transformed quality levels: [High, good, fair, poor, low] and $\mathcal{P}$ represents the model’s output logit of each quality level. Specifically, we first extract the vector (with its dimension matching the tokenizer vocabulary size) at the model’s output sequence corresponding to the quality-level description word (the $-3$ index in our model). We then take the logits at the indexes in this vector associated with each quality level in the tokenizer’s vocabulary (indexes $1550$, $1661$, $6624$, $7852$, and $3347$ in our model). Then, we apply softmax normalization to these $5$ logits. $\omega$ denotes the weights assigned to the normalized probabilities of these $5$ quality levels ($[1,0.75,0.5,0.25,0]$). The weighted sum of these probabilities yields the predicted quality score $\mathcal{Q}$, ranging between $[0,1]$.
@@
Figure 9: A video with extremely high color temperature, very low sharpness, and frequent, severe stuttering. The case demonstrates that the Assistant model can accurately identify videos with low fluency, perceive the level of color temperature, give reasonable suggestions for quality refinement, and provide a relatively accurate scoring for low-quality videos.
Figure 10: A video shot on a mobile phone with a stable camera trajectory. The video undergoes some compression during file transfer, slightly reducing its sharpness. At the beginning, the video shows noticeable overexposure on the window, which does not persist throughout, resulting in overall medium quality. This example demonstrates that the Assistant model can accurately detect camera shake, perform precise spatial-temporal retrieval or analysis of distortions, and provide a reasonably accurate rating for a medium-quality video.
@@ Appendix F Applications
We created a demo based on Gradio to test the application of the Assistant model in video quality question-answering. We provide four specific examples in Fig. 9, 10, 11, and 12) .
@@ Appendix G Limitations
Due to time constraints, we can not conduct extensive experiments to obtain a larger amount of distortion recognition pretraining data. We utilize the data from existing studies and some synthetically generated data. Therefore, the Stage-1 can be regarded as a ”task-specific pretraining for video quality assessment” on small-scale data rather than the large-scale pretraining typically used to construct foundation models in comprehensive LMMs.
@@
Despite employing various methods, such as GPT-based extension and refinement and quality-level-referenced attribute selection to minimize annotation biases, the individual perception biases are impossible to eliminate entirely. This limitation slightly impacts the diversity of model outputs and the accuracy of responses to specific questions.
Furthermore, since the series models (especially the Assistant) are specifically designed for low-level video quality assessment tasks, especially for the UGC videos, their performance and generalization ability might face a little challenge on other general-purpose tasks/benchmarks and video content types.
@@ Appendix H Acknowledgements
In the subjective annotation experiments, all participants are clearly informed in advance about the workload. Participation in the experiment is entirely voluntary without any enforcement. Throughout the experiment, no participants report experiencing fatigue or discomfort. Upon completion of the experiment, we provide all participants with appropriate labor wages. We would like to express our heartfelt gratitude to all participants for their contributions to the subjective experiments.
@@ Appendix I License
All the outputs of our work, including the complete data of the VQA2 Instruction Dataset as well as the training and evaluation code and fine-tuned weights of the VQA2 Series Models, will be fully open-sourced and made freely available to the research community for further studies.